OpenStudy (tmagloire1):

Write equation is slope intercept. The line parallel to y=-7x-2 through (-4,6)

OpenStudy (anonymous):

y-y1=m(x-x1) subsitiue in y-(-4)= m(x-6) y+4= m (x-6)

OpenStudy (anonymous):

did u find the slope already?

OpenStudy (tmagloire1):

Oh yeah I forgot about point slope. Thanks

OpenStudy (tmagloire1):

uhm 1 sec im finding it now

OpenStudy (anonymous):

u welcome